Session: May 2001 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Paco Colomer at Thu May 31 11:02:55 METDST 2001) Minor failures
Good weather. No major problems, but 30 minutes of data lost due to problems with the antenna control computer.

Medicina (orfei at Thu May 31 11:35:25 METDST 2001) Success
No problems to comment

Onsala (A. Polatidis at Thu May 31 11:46:12 METDST 2001) Success

Noto (P. Leto at Thu May 31 13:11:52 METDST 2001) Success
No known problem.

Effelsberg (A. Kraus at Thu May 31 17:00:40 METDST 2001) Success

Shanghai (Huang Xinyong at Fri Jun 1 02:47:55 METDST 2001) Success
No know problems.

Westerbork (Tony Foley at Mon Jun 4 11:39:12 METDST 2001) Minor failures
Lost 05:17-06:02 with array problem. Bad weather for 3.6cm - heavy rain showers with some thunder.

Urumqi (Liu Xiang at Sun Jun 24 13:54:39 METDST 2001) Minor failures
01h00m---01h46m tape does not move. vc1,3,5,7 parity is high.